# Mordell–Weil groups over large algebraic extensions of fields of characteristic zero

### v1

@@ 1. Introduction
In this paper, we study the structure of the Mordell–Weil groups of semiabelian varieties over large algebraic extensions of a finitely generated field of characteristic zero. In 1974, Frey and Jarden [FreJ74, p. 127] conjectured that, if $A$ is an abelian variety over an algebraic number field $k$, then the Mordell–Weil group $A(k^{\mathrm{ab}})$ of $A$ over the the maximal abelian extension $k^{\mathrm{ab}}$ of $k$ is of infinite rank. There have been many works related to this conjecture ([RW02, Im06, Kob06, Pet06, SY12, IL13], to cite a few). Some authors [Moo09, Moo12, GHP15] have proved that, for certain algebraic extensions $K/k$ of infinite degree, the Mordell–Weil group $A(K)$ is not only of infinite rank but also free modulo torsion. In particular, $A(K)$ modulo torsion has trivial divisible part; it is this property that we focus on in this paper. A perfect field $K$ is called Kummer-faithful if, for any semiabelian variety $A$ over a finite extension $L$ of $K$, the Mordell-Weil group $A(L)$ has trivial divisible part (for more detail, see Definition 1.1). In this paper, we obtain a wide class of Kummer-faithful fields that are not sub-$p$-adic (see Definition 1.1 for the definition). This improves some of the previous results by Ozeki–Taguchi [OT22] and by Ohtani [Oht22] (see also its corrigendum [Oht23]). The notion of Kummer-faithfulness (resp. sub-$p$-adicness) originates from anabelian geometry and it is expected (resp. known) that a field possessing this property makes it suitable as a ground field for anabelian geometry; for more detail, see, e.g., [Moc99, Moc15, Hos17].
@@
[…]
@@
Note that, if $m_{p}\geq 1$ for infinitely many primes $p$, then the field $K_{g,\boldsymbol{m}}$ is not sub-$p$-adic, since the smaller field $k_{g,\boldsymbol{m}}$ is already not sub-$p$-adic [OT22, Remark 3.4].
@@
We also investigate the Kummer-faithfulness of $\overline{K}(\sigma)$ and of $\overline{K}[\sigma]$. Ohtani showed that, if $K$ is a number field and $e\geq 2$, then any finite extension of $\overline{K}[\sigma]$ is Kummer-faithful for almost all $\sigma\in G_{K}^{e}$ (see [Oht22, Corollary 1] and its corrigendum [Oht23, Corollary 1]). Our next result is an extension of the result by Ohtani and gives a nearly complete answer for the Kummer-faithfulness for finite extensions of $\overline{K}(\sigma)$ and of $\overline{K}[\sigma]$ when $K$ is a finitely generated field over $\mathbb{Q}$. We remark that neither $\overline{K}(\sigma)$ nor $\overline{K}[\sigma]$ are sub-$p$-adic for almost all $\sigma\in G_{K}^{e}$ and any prime number $p$ (see Proposition 5.1).
@@
    ($=$ Theorem 5.3) Assume $e\geq 1$. Any finite extension of $\overline{K}[\sigma]$ is Kummer-faithful for almost all $\sigma\in G_{K}^{e}$.
@@
    ($=$ Theorem 5.2) Assume $e\geq 2$. Any finite extension of $\overline{K}(\sigma)$ is Kummer-faithful for almost all $\sigma\in G_{K}^{e}$.
  3. (3)
    ($=$ Theorem 5.4) Any finite extension of $\overline{K}(\sigma)$ is torally Kummer-faithful for almost all $\sigma\in G_{K}$.
We note that the absolute Galois group $G_{\overline{K}(\sigma)}$ of $\overline{K}(\sigma)$ is the closed subgroup $\langle\sigma\rangle$ in $G_{K}$ generated by the entries of $\sigma\in G_{K}^{e}$ and it is abelian if $e=1$. Theorem 1.5 (3) answers in the negative the questions [MT22, Remark 2.4.1, Questions 1 and 2] posed by Minamide–Tsujimura which asked whether the absolute Galois group of any torally Kummer-faithful field is slim (i.e., every open subgroup has trivial center).
@@
Section 4 and a part of Section 5 are originally part of the PhD thesis [Asa24] of the first author.
@@ Acknowledgments
The authors thank Naganori Yamaguchi for helpful discussions during the preparation of this paper. They are grateful to Benjamin Collas, Shinichi Mochizuki, Yoshiyasu Ozeki, Akio Tamagawa, and Shota Tsujimura for their useful comments on an early version of the manuscript. This work was supported by JSPS KAKENHI Grant Number JP23K03068 and by the Research Institute for Mathematical Sciences, an International Joint Usage/Research Center located in Kyoto University.
@@
We begin by proving two lemmas needed in the proof of Theorem 1.2 ($=$ Theorem 2.3). Let $K$ be a complete discrete valuation field with perfect residue field. If $L/K$ is a finite Galois extension, then the Galois group $G$ has two filtrations; the lower-numbering filtration $(G_{u})_{u\in\mathbb{R}}$ and the upper-numbering filtration $(G^{v})_{v\in\mathbb{R}}$. These are related by the Herbrand function $\varphi_{L/K}$ defined by
@@
Assume $K$ has finite residue field with absolute ramification index $e_{K}$ and residue characteristic $p_{K}$. If $L/K$ is a finite abelian extension of exponent $p^{m}$ for some prime number $p$ and integer $m\geq 1$, then it has ramification bounded by
@@
If $K$ has characteristic $>0$, then we understand that $e_{K}=\infty$, so that the lemma trivially holds true.
@@
If $p_{K}\not=p$, then $L/K$ is tamely ramified and hence it has ramification bounded by $1$. Suppose $K$ has residue characteristic $p$. We may assume $L/K$ is totally ramified and cyclic of degree $p^{m}$. Let ${\mathcal{O}}_{K}$ denote the valuation ring of $K$. By local class field theory, we have the reciprocity map
@@
so that any cyclic quotient of $\mathrm{Gal}(L/K)$ of degree dividing $p^{m}$ has maximal ramification break $\leq u$. This follow from the equality
$1+x\ =\ \left(\exp\left(\frac{1}{p^{m}}\log(1+x)\right)\right)^{p^{m}},$
@@
This follows from the following lemma, since if $P\in A(k)$ is a non-torsion point, then it is divisible by $p^{m_{p}}$ ($>1$) in $A(K_{g,\boldsymbol{m}})$ for infinitely many $p$.
@@ 3. Known results for our second type of extensions
      Assume $e=1$. For almost all $\sigma\in G_{K}$ and any positive integer $d$, there exist infinitely many prime numbers $\ell$ such that $[\overline{K}(\sigma)(\zeta_{\ell}):\overline{K}(\sigma)]=d$. In particular, for almost all $\sigma\in G_{K}$ and any finite extension $M$ of $\overline{K}(\sigma)$, there exist infinitely many roots of unity contained in $M$ and a prime number $\ell$ such that $\zeta_{\ell}\notin M$.
@@
      Assume $e\geq 2$. Then for almost all $\sigma\in G_{K}^{e}$ and any positive integer $d$, there exist only finitely many positive integers $n$ not divisible by the characteristic of $K$ such that $[\overline{K}(\sigma)(\zeta_{n}):\overline{K}(\sigma)]\leq d$. In particular, for almost all $\sigma\in G_{K}^{e}$ and any finite extension $M$ of $\overline{K}(\sigma)$, there are only finitely many roots of unity contained in $M$.
@@
Larsen [Lar03] conjectured that the statement on the field $\overline{K}(\sigma)$ in (1) holds for any $\sigma\in G_{K}^{e}$. He and Im [IL08, Theorem 1.4] gave an affirmative answer to this conjecture under the assumptions that $K$ is of characteristic different from two and $e=1$. Geyer–Jarden [GJ78] conjectured that Statements (a)–(c) in (3) in this theorem hold for any finitely generated field $K$ over its prime field. We note that the paper of Jacobson–Jarden [JJ84] involves a proof of Statement (a) for $K$ with positive characteristic, but it contains an error as indicated in [JJ85]. Statements (a) and (b) for $K$ which is infinite and has positive characteristic remain open.
@@
Let $G$ be a profinite group with the normalized Haar measure $\mu_{G}$. A family $\{B_{i}\}_{i\in I}$ of measurable subsets of $G$ is $\mu_{G}$-independent if $\mu_{G}(\bigcap_{i\in J}B_{i})=\prod_{i\in J}\mu_{G}(B_{i})$ for every finite subset $J$ in $I$.
Our theorems involving the normalized Haar measure are proved using the following well-known lemma in probability theory. We note that the proof via this lemma is non-constructive; it does not give an explicit element for which our statement holds. We also notice that the proofs of some of our theorems are not explicitly described in their use of this lemma; most of the assertions in Theorem 3.1 rely on this lemma and our theorems ultimately depend on it through such assertions.
Lemma 3.4 (Borel–Cantelli lemma).
Let $\{B_{i}\}_{i=1}^{\infty}$ be a countable collection of measurable subsets of a profinite group $G$ with the normalized Haar measure $\mu_{G}$. Define
[…]
  1. (1)
    (First Borel–Cantelli lemma) If $\sum_{i=1}^{\infty}\mu_{G}(B_{i})<\infty$, then $\mu_{G}(B)=0$.
  2. (2)
    (Second Borel–Cantelli lemma) Suppose that $\{B_{i}\}_{i=1}^{\infty}$ is $\mu_{G}$-independent. If $\sum_{i=1}^{\infty}\mu_{G}(B_{i})=\infty$, then $\mu_{G}(B)=1$.
Proof.
See [FriJ23, Lemma 21.3.5]. ∎
@@ 5. Kummer-faithfulness for some large algebraic extensions
In the present section, we determine when the fields $\overline{K}(\sigma)$ and $\overline{K}[\sigma]$ are Kummer-faithful for a finitely generated field $K$ over $\mathbb{Q}$ and $\sigma\in G_{K}^{e}$. The following proposition suggests that these fields should be of interest since every sub-$p$-adic field is Kummer-faithful.
@@
Since $\overline{K}(\sigma)$ contains $\overline{K}[\sigma]$, we only have to discuss whether $\overline{K}[\sigma]$ is sub-$p$-adic or not. We first consider the case where $K=k$ is a number field. It is known that $\overline{k}[\sigma]$ is pseudo algebraically closed (see [FriJ23, Chapter 12] for the definition) for almost all $\sigma\in G_{k}^{e}$ [FriJ23, Theorem 21.12.2]. Moreover, an algebraic extension of a pseudo algebraically closed field is henselian with respect to some nonarchimedean place if and only if it is the algebraic closure [FriJ23, Corollary 12.5.6]. Then we have $\overline{k}\cap(\overline{k}[\sigma])_{\mathfrak{p}}=\overline{k}$ for almost all $\sigma\in G_{k}^{e}$ and any nonarchimedean place $\mathfrak{p}$ of $\overline{k}[\sigma]$. If $\overline{k}[\sigma]$ is sub-$p$-adic, then it is a subfield of a finite extension of $\mathbb{Q}_{p}$. However, since ¯$\overline{\mathbb{Q}}\cap L\neq\overline{\mathbb{Q}}$ for any finite extension $L$ of $\mathbb{Q}_{p}$, the field $\overline{k}[\sigma]$ is not sub-$p$-adic for almost all $\sigma\in G_{k}^{e}$.
@@
Theorem 5.2.
Suppose that $K$ is a finitely generated field over $\mathbb{Q}$ and $e\geq 2$. Then, for almost all $\sigma\in G_{K}^{e}$, any finite extension of $\overline{K}(\sigma)$ is Kummer-faithful.
Proof.
Since Kummer-faithfulness is preserved under finite extensions, we only have to show that $\overline{K}(\sigma)$ is Kummer-faithful for almost all $\sigma\in G_{K}^{e}$. We know from Theorem 3.1 (4) that $A(\overline{K}(\sigma))_{\mathrm{div}}=0$ for almost all $\sigma\in G_{K}^{e}$ and any abelian variety $A$ over $\overline{K}(\sigma)$. By Proposition 3.3, it suffices to show $\mathbb{G}_{\mathrm{m}}(M)_{\mathrm{div}}=0$ for almost all $\sigma\in G_{K}^{e}$ and any finite extension $M$ of $\overline{K}(\sigma)$.
Let $S$ be the set of $\sigma\in G_{K}^{e}$ such that $\mathbb{G}_{\mathrm{m}}(M)_{\mathrm{div}}\neq 0$ for some finite extension $M$ of $\overline{K}(\sigma)$. For every $a\in\overline{K}^{\times}\smallsetminus\{1\}$, let $S_{a}$ be the set of $\sigma\in G_{K}^{e}$ such that $a\in\mathbb{G}_{\mathrm{m}}(M)_{\mathrm{div}}$ for some finite extension $M$ of $\overline{K}(\sigma)$. Then we have $S=\bigcup_{a\in\overline{K}^{\times}\smallsetminus\{1\}}S_{a}$. Since $\overline{K}^{\times}\smallsetminus\{1\}$ is countable, it suffices to show $\mu(S_{a})=0$ for each $a\in\overline{K}^{\times}\smallsetminus\{1\}$.
If $a$ is a root of unity, then some finite extension of $\overline{K}(\sigma)$ has all roots of unity for each $\sigma\in S_{a}$. Since the set of $\sigma\in G_{K}^{e}$ such that some finite extension of $\overline{K}(\sigma)$ contains infinitely many roots of unity has measure zero with respect to $\mu$ by Theorem 3.1 (2-ii), we have $\mu(S_{a})=0$.
Suppose that $a$ is not a root of unity. For $n\geq 1$, put
$T_{a}^{(n)}=\{\sigma\in G_{K}^{e}\mid\text{some $n$-th root of $a$ belongs to }\overline{K}(\sigma)\},$
and let $\alpha^{(n)}_{1},\ldots,\alpha^{(n)}_{n}$ be all solutions of $X^{n}=a$ in $\overline{K}$. Let $T_{a}^{(n,i)}$ be the set of $\sigma\in G_{K}^{e}$ such that $\alpha^{(n)}_{i}\in\overline{K}(\sigma)$. Then we have $T_{a}^{(n)}=\bigcup_{i=1}^{n}T_{a}^{(n,i)}$ and $\mu(T_{a}^{(n,i)})=1/[K(\alpha^{(n)}_{i}):K]^{e}$. The assumption on $a$ tells us that there exists a prime number $\ell_{0}$ such that $a$ is not an $\ell$-th power in $K(a)$ for all prime numbers $\ell\geq\ell_{0}$. If $\ell$ is such a prime, then the polynomial $X^{\ell}-a$ is irreducible in $K(a)[X]$ [Lan02, Chapter VI, Theorem 9.1]. Hence we have
$\Big{[}K(\alpha^{(\ell)}_{i}):K\Big{]}=\Big{[}K(\alpha^{(\ell)}_{i}):K(a)\Big{]}[K(a):K]\geq\ell$
and
$\mu\left(T_{a}^{(\ell)}\right)=\mu\left(\bigcup_{i=1}^{\ell}T_{a}^{(\ell,i)}\right)\leq\sum_{i=1}^{\ell}\mu\left(T_{a}^{(\ell,i)}\right)\leq\frac{\ell}{\ell^{e}}=\frac{1}{\ell^{e-1}}.$
Since $e\geq 2$, we can take an infinite set $\Lambda_{a}$ of prime numbers $\ell\geq\ell_{0}$ with $\sum_{\ell\in\Lambda_{a}}1/\ell^{e-1}<\infty$. The first Borel–Cantelli lemma (Lemma 3.4 (1)) implies that, for almost all $\sigma\in G_{K}^{e}$, there are only finitely many $\ell\in\Lambda_{a}$ with $\sigma\in T_{a}^{(\ell)}$. This shows that there are infinitely many $\ell\in\Lambda_{a}$ such that $a$ is not an $\ell$-th power in $\overline{K}(\sigma)$ for such $\sigma$. By the same argument as above, we obtain
$\Big{[}\overline{K}(\sigma)(\alpha^{(\ell)}_{i}):\overline{K}(\sigma)\Big{]}\geq\ell$
for infinitely many $\ell\in\Lambda_{a}$ and all $1\leq i\leq\ell$. This implies that $a\notin\mathbb{G}_{\mathrm{m}}(M)_{\mathrm{div}}$ for almost all $\sigma\in G_{K}^{e}$ and all finite extensions $M$ of $\overline{K}(\sigma)$, which yields $\mu(S_{a})=0$, as desired. ∎
@@
If $e\geq 2$, then the proposition follows from Theorem 5.2. (We remark that it also follows from Theorems 3.1 (4) and 4.1.) Hence we may assume $e=1$. Since Kummer-faithfulness is preserved under finite extensions, we only discuss whether $\overline{K}[\sigma]$ is Kummer-faithful. By Proposition 3.3, it suffices to show that the following two statements hold for almost all $\sigma\in G_{K}$:
@@
    $\mathbb{G}_{\mathrm{m}}(L)_{\mathrm{div}}=0$ for any finite extension $L$ of $\overline{K}[\sigma]$.
@@
In the remainder of this paper, we discuss the Kummer-faithfulness for the field $\overline{K}(\sigma)$ in the case where $e=1$. By Proposition 3.3, we can divide the discussion into two parts; torally Kummer-faithfulness and the vanishing property of the divisible part of the Mordell–Weil groups of abelian varieties. Jarden–Petersen [JP22, Remark 5.5], who proved the latter part when $e\geq 2$ (cf. Theorem 3.1 (4)), pointed out that their proof does not work when $e=1$. We do not know at the time of writing this paper whether the latter part establishes if $e=1$. However, we demonstrate that the former part still holds even in this case.
@@
Suppose that $K$ is a finitely generated field over $\mathbb{Q}$. Then, for almost all $\sigma\in G_{K}$, any finite extension of $\overline{K}(\sigma)$ is torally Kummer-faithful.
@@
If we accepted the claim in [Moc15, Theorem 1.11] that the absolute Galois group of any Kummer-faithful field of characteristic zero is slim, then the field $\overline{K}(\sigma)$ would not be Kummer-faithful for any $\sigma\in G_{K}$ since the absolute Galois group $G_{\overline{K}(\sigma)}$ of $\overline{K}(\sigma)$ is abelian. However, Mochizuki recently informed us that there is a gap in the proof of this claim and the status remains open for general Kummer-faithful fields of characteristic zero. If the Kummer-faithfulness result for $\overline{K}(\sigma)$ holds when $e=1$, then it leads us to conclude that the above claim does not hold in the general case, although it provides no explicit counterexample.
@@
Before proving Theorem 5.4, we need a lemma concerning the degree of the splitting field of the polynomial $X^{n}-a$ with some conditions on $a\in K$ and a positive integer $n$.
A family $\{L_{i}\}_{i\in I}$ of finite separable extensions of a field $K$ is linearly disjoint over $K$ if $[K(\bigcup_{i\in J}L_{i}):K]=\prod_{i\in J}[L_{i}:K]$ for every finite subset $J$ in $I$. This occurs if and only if the family $\{G_{L_{i}}\}_{i\in I}$ of open subgroups of $G_{K}$ is $\mu_{G_{K}}$-independent [FriJ23, Lemma 21.5.1].
@@
  • if $n=\ell_{1}\cdots\ell_{r}$, where $\ell_{1},\ldots,\ell_{r}$ are distinct prime numbers in $\Lambda_{a}$, then the splitting field of $X^{n}-a$ over $K$ is of degree $n\varphi(n)$ over $K$, where $\varphi$ denotes the Euler totient function.

### v2

@@ 1. Introduction
In this paper, we study the structure of the Mordell–Weil groups of semiabelian varieties over large algebraic extensions of a finitely generated field of characteristic zero. In 1974, Frey and Jarden [FreJ74, p. 127] conjectured that, if $A$ is an abelian variety over an algebraic number field $k$, then the Mordell–Weil group $A(k^{\mathrm{ab}})$ of $A$ over the the maximal abelian extension $k^{\mathrm{ab}}$ of $k$ is of infinite rank. There have been many works related to this conjecture ([RW02, Im06, Kob06, Pet06, SY12, IL13a], to cite a few). Some authors [Moo09, Moo12, GHP15] have proved that, for certain algebraic extensions $K/k$ of infinite degree, the Mordell–Weil group $A(K)$ is not only of infinite rank but also free modulo torsion. In particular, $A(K)$ modulo torsion has trivial divisible part; it is this property that we focus on in this paper. A perfect field $K$ is called Kummer-faithful if, for any semiabelian variety $A$ over a finite extension $L$ of $K$, the Mordell–Weil group $A(L)$ has trivial divisible part (for more detail, see Definition 1.1). In this paper, we obtain a wide class of Kummer-faithful fields that are not sub-$p$-adic (see Definition 1.1 for the definition). This improves some of the previous results by Ozeki–Taguchi [OT22] and by Ohtani [Oht22] (see also its corrigendum [Oht23]). The notion of Kummer-faithfulness (resp. sub-$p$-adicness) originates from anabelian geometry and it is expected (resp. known) that a field possessing this property makes it suitable as a ground field for anabelian geometry.
@@
    for any prime number $\ell\not=\text{\rm{ch}}(k)$ and any $i,r$ with $i\not=2r$, where the suffix $(\text{--})_{G_{L}}$ means the largest quotient of $(\text{--})$ on which the absolute Galois group $G_{L}$ of $L$ acts trivially.
@@
Recall also that the Grothendieck conjecture has been proved for hyperbolic curves over sub-$p$-adic fields [Moc99] and, later, in [Moc15], [Hos17] etc., Mochizuki and Hoshi try to extend such results to hyperbolic curves over Kummer-faithful fields. Thus it would be interesting to supply many examples of Kummer-faithful fields that are not sub-$p$-adic.
@@
This extends Theorem 3.3 of [OT22], which proved the same for the field $k_{g,\boldsymbol{m}}$, to the case of much larger field $K_{g,\boldsymbol{m}}$. Note that, if $m_{p}\geq 1$ for infinitely many primes $p$, then the field $K_{g,\boldsymbol{m}}$ is not sub-$p$-adic, since the smaller field $k_{g,\boldsymbol{m}}$ is already not sub-$p$-adic [OT22, Remark 3.4].
@@
We also investigate the Kummer-faithfulness of $\overline{K}(\sigma)$ and of $\overline{K}[\sigma]$. Ohtani showed that, if $K$ is a number field and $e\geq 2$, then any finite extension of $\overline{K}[\sigma]$ is Kummer-faithful for almost all $\sigma\in G_{K}^{e}$ (see [Oht22, Corollary 1] and its corrigendum [Oht23, Corollary 1]). Our next result contains an improvement over the one by Ohtani, that is, we show that the same conclusion holds even if $e=1$ for arbitrary finitely generated field $K$ over $\mathbb{Q}$. For the field $\overline{K}(\sigma)$, we obtain a partial result on its multiplicative group for any $e\geq 1$. We remark that neither $\overline{K}(\sigma)$ nor $\overline{K}[\sigma]$ are sub-$p$-adic for almost all $\sigma\in G_{K}^{e}$ and any prime number $p$ (see Proposition 5.1).
@@
    (== Theorem 5.2) Any finite extension of $\overline{K}[\sigma]$ is Kummer-faithful for almost all $\sigma\in G_{K}^{e}$.
@@
    (== Theorem 5.3) The multiplicative group of $\overline{K}(\sigma)$ has no nontrivial divisible point for almost all $\sigma\in G_{K}^{e}$.
@@
Section 4 and a part of Section 5 are originally part of the PhD thesis [Asa24] of the first author. Theorem 4.2.1 in the thesis, which claimed the (torally) Kummer-faithfulness property for $\overline{K}(\sigma)$, is excluded in this paper because its proof was incorrect. Note that Theorem 1.5 (2) (== Theorem 5.3) in this paper is the corrected (weaker) version of this assertion.
@@ Acknowledgments
The authors thank Naganori Yamaguchi for helpful discussions during the preparation of this paper. They are grateful to Benjamin Collas, Shinichi Mochizuki, Yoshiyasu Ozeki, Akio Tamagawa, and Shota Tsujimura for their useful comments on an early version of the manuscript. This work was supported by the Research Institute for Mathematical Sciences, an International Joint Usage/Research Center located in Kyoto University. The second author was supported by JSPS KAKENHI Grant Number JP23K03068.
This paper has been published in Res. Number Theory 11 (2025), Paper No. 89, 18 pp. (https://doi.org/10.1007/s40993-025-00658-2) Open Access funding provided by Institute of Science Tokyo.
@@
We begin by proving two lemmas needed in the proof of Theorem 1.2 (== Theorem 2.3). Let $K$ be a complete discrete valuation field with perfect residue field. If $L/K$ is a finite Galois extension, then the Galois group $G$ has two filtrations; the lower-numbering filtration $(G_{u})_{u\in\mathbb{R}}$ and the upper-numbering filtration $(G^{v})_{v\in\mathbb{R}}$. These are related by the Herbrand function $\varphi_{L/K}$ defined by
@@
Assume that $K$ has finite residue field with absolute ramification index $e_{K}$ and residue characteristic $p_{K}$. If $L/K$ is a finite abelian extension of exponent $p^{m}$ for some prime number $p$ and integer $m\geq 1$, then it has ramification bounded by
@@
If $K$ has characteristic >0>0, then we understand that $e_{K}=\infty$, so that the lemma trivially holds true.
@@
If $p_{K}\not=p$, then $L/K$ is tamely ramified and hence it has ramification bounded by $1$. Suppose that $K$ has residue characteristic $p$. We may assume that $L/K$ is totally ramified and cyclic of degree $p^{m}$. Let ${\mathcal{O}}_{K}$ denote the valuation ring of $K$. By local class field theory, we have the reciprocity map
@@
so that any cyclic quotient of $\mathrm{Gal}(L/K)$ of degree dividing $p^{m}$ has maximal ramification break $\leq u$. This follows from the equality
[…]
@@
This follows from the following lemma, since if $P\in A(k)$ is a non-torsion point, then it is divisible by $p^{m_{p}}$ (>1>1) in $A(K_{g,\boldsymbol{m}})$ for infinitely many $p$.
@@ 3. Known results for our second type of extensions
      Assume $e=1$. For almost all $\sigma\in G_{K}$ and any positive integer $d$, there exist infinitely many prime numbers $\ell$ such that $\overline{K}(\sigma)(\zeta_{\ell}):\overline{K}(\sigma)]=d$. In particular, for almost all $\sigma\in G_{K}$ and any finite extension $M$ of $\overline{K}(\sigma)$, there exist infinitely many roots of unity contained in $M$ and a prime number $\ell$ such that $\zeta_{\ell}\notin M$.
@@
      Assume $e\geq 2$. Then for almost all $\sigma\in G_{K}^{e}$ and any positive integer $d$, there exist only finitely many positive integers $n$ not divisible by the characteristic of $K$ such that $\overline{K}(\sigma)(\zeta_{n}):\overline{K}(\sigma)]\leq d$. In particular, for almost all $\sigma\in G_{K}^{e}$ and any finite extension $M$ of $\overline{K}(\sigma)$, there are only finitely many roots of unity contained in $M$.
@@
The results, including ours, involving the normalized Haar measure are non-constructive as they are proved by calculating the measures of various subsets in $G_{K}^{e}$. Thus they do not give an explicit element for which their statement holds. Larsen [Lar03] conjectured that the statement on the field $\overline{K}(\sigma)$ in (1) holds for any $\sigma\in G_{K}^{e}$. He and Im [IL08, Theorem 1.4] gave an affirmative answer to this conjecture under the assumptions that $K$ is of characteristic different from two and $e=1$. Im with collaborators [BI08, IL13b, CI25] has obtained other results concerning Larsen’s conjecture. See also the survey [IL21] by Im and Larsen.
Geyer and Jarden [GJ78] conjectured that Statements (a)–(c) in (3) in this theorem hold for any finitely generated field $K$ over its prime field. We note that the paper of Jacobson and Jarden [JJ84] involves a proof of Statement (a) for $K$ with positive characteristic, but it contains an error as indicated in [JJ85]. Statements (a) and (b) for $K$ which is infinite and has positive characteristic remain open.
@@ 5. Kummer-faithfulness for some large algebraic extensions
In the present section, we investigate when the fields $\overline{K}(\sigma)$ and $\overline{K}[\sigma]$ are Kummer-faithful for a finitely generated field $K$ over $\mathbb{Q}$ and $\sigma\in G_{K}^{e}$. The following proposition suggests that these fields should be of interest since every sub-$p$-adic field is Kummer-faithful.
@@
Since $\overline{K}(\sigma)$ contains $\overline{K}[\sigma]$, we only have to discuss whether $\overline{K}[\sigma]$ is sub-$p$-adic or not. We first consider the case where $K=k$ is a number field. It is known that $\overline{k}[\sigma]$ is pseudo algebraically closed (see [FriJ23, Chapter 12] for the definition) for almost all $\sigma\in G_{k}^{e}$ [FriJ23, Theorem 21.12.2]. Moreover, an algebraic extension of a pseudo algebraically closed field is henselian with respect to some nonarchimedean place if and only if it is the algebraic closure [FriJ23, Corollary 12.5.6]. Then we have $\overline{k}\cap(\overline{k}[\sigma])_{\mathfrak{p}}=\overline{k}$ for almost all $\sigma\in G_{k}^{e}$ and any nonarchimedean place $\mathfrak{p}$ of $\overline{k}[\sigma]$. If $\overline{k}[\sigma]$ is sub-$p$-adic, then it is a subfield of a finite extension of $\mathbb{Q}_{p}$. However, since $\overline{\mathbb{Q}}\cap L\neq\overline{\mathbb{Q}}$ for any finite extension $L$ of $\mathbb{Q}_{p}$, the field $\overline{k}[\sigma]$ is not sub-$p$-adic for almost all $\sigma\in G_{k}^{e}$.
@@
If $e\geq 2$, then the proposition follows from Corollary 4.3. Hence we may assume $e=1$. Since Kummer-faithfulness is preserved under finite extensions, we only discuss whether $\overline{K}[\sigma]$ is Kummer-faithful. By Proposition 3.3, it suffices to show that the following two statements hold for almost all $\sigma\in G_{K}$:
@@
    $\mathbb{G}_{\mathrm{m}}(L)_{\mathrm{div}}=0$ for any finite extension $L$ of $\overline{K}[\sigma]$;
@@
In the remainder of this paper, we discuss the Kummer-faithfulness for the field $\overline{K}(\sigma)$. By Proposition 3.3, we can divide the discussion into two parts; torally Kummer-faithfulness and the vanishing property of the divisible part of the Mordell–Weil groups of abelian varieties. Jarden and Petersen [JP22, Remark 5.5], who proved the latter part when $e\geq 2$ (cf. Theorem 3.1 (4)), pointed out that their proof does not work when $e=1$. We do not know at the time of writing this paper whether the Kummer-faithfulness property holds for any $e\geq 1$. Here we give a partial result on the torally Kummer-faithfulness for any $e\geq 1$.
@@
Suppose that $K$ is a finitely generated field over $\mathbb{Q}$ and $e$ is a positive integer. Then, the group $\mathbb{G}_{\mathrm{m}}(\overline{K}(\sigma))_{\mathrm{div}}$ is trivial for almost all $\sigma\in G_{K}^{e}$.
@@
If we accepted the claim in [Moc15, Theorem 1.11] that the absolute Galois group of any Kummer-faithful field of characteristic zero is slim (i.e., every open subgroup has trivial center), then the field $\overline{K}(\sigma)$ would not be Kummer-faithful for any $\sigma\in G_{K}$ since the absolute Galois group $G_{\overline{K}(\sigma)}$ of $\overline{K}(\sigma)$ is abelian; $G_{\overline{K}(\sigma)}$ is the closed subgroup $\langle\sigma\rangle$ in $G_{K}$ generated by $\sigma$. However, Mochizuki recently informed us that there is a gap in the proof of this claim and the status remains open for general Kummer-faithful fields of characteristic zero. If the Kummer-faithfulness result for $\overline{K}(\sigma)$ holds when $e=1$, then it leads us to conclude that the above claim does not hold in the general case, although it provides no explicit counterexample. It would also answer in the negative the questions [MT22, Remark 2.4.1, Questions 1 and 2] posed by Minamide and Tsujimura, which ask whether the absolute Galois group of any torally Kummer-faithful field is slim.
@@
It is obvious that we may assume $e=1$ to prove Theorem 5.3. Before proving the theorem, we need a lemma concerning the degree of the splitting field of the polynomial $X^{n}-a$ with some conditions on $a\in K$ and a positive integer $n$.
A family $\{B_{i}\}_{i\in I}$ of measurable subsets of $G_{K}$ is $\mu$-independent if $\mu(\bigcap_{i\in J}B_{i})=\prod_{i\in J}\mu(B_{i})$ for every finite subset $J$ in $I$. A family $\{L_{i}\}_{i\in I}$ of finite separable extensions of a field $K$ is linearly disjoint over $K$ if $K(\bigcup_{i\in J}L_{i}):K]=\prod_{i\in J}[L_{i}:K]$ for every finite subset $J$ in $I$. This occurs if and only if the family $\{G_{L_{i}}\}_{i\in I}$ of open subgroups of $G_{K}$ is $\mu$-independent [FriJ23, Lemma 21.5.1].
@@
Let $\varphi$ denote the Euler totient function, i.e., $\varphi(n)$ for a positive integer $n$ is the number of positive integers less than or equal to $n$ that are coprime to $n$.
@@
  • if $n=\ell_{1}\cdots\ell_{r}$, where $\ell_{1},\ldots,\ell_{r}$ are distinct prime numbers in $\Lambda_{a}$, then the splitting field of $X^{n}-a$ over $K$ is of degree $n\varphi(n)$ over $K$.